OpenStudy (anonymous):

Check my answer please. If i got it right please dont explain how to do it. If its wrong please explain why and how to get it right :) Find the product of (x + 3)^2. I got x^2 + 9

OpenStudy (haseeb96):

no u are wrong just one will be added in yur answer that this is the formula (a+b)^2 = a^2 +2ab + b^2 (x+3)^2 = x^2 + 2(x)(3) +9 = x^2 +6x + 9 is the correct answer

OpenStudy (anonymous):

ohh, i get it. Thank you

OpenStudy (haseeb96):

for (a-b)^2 = a^2 -2ab + b^2 for yur extra knowledge because there must be this type of questions will come in yur test or paper or in yur homework